OpenStudy (anonymous):

write the equation -3x+2y=7 in slope-intercept form

pooja195 (pooja195):

Add 3x to both sides 2y=3x+7 divide all of the numbers by 2

pooja195 (pooja195):

^to isolate y

OpenStudy (anonymous):

so its y=-3/2x+7

pooja195 (pooja195):

Divide the 7 too...

OpenStudy (anonymous):

so its y=-3/2x+7/2

pooja195 (pooja195):

y=-3/2x ^is correct so far but you need to put the 7/2 because we are dividing all numbers by 2 y=-3/2x+7/2 <----Correct :)